factor trinomial below:
-7x^2 + 5x + 12 = 0

Answer by TimothyLamb(4379)

-7xx + 5x + 12 = 0
---
the above quadratic equation is in standard form, with a=-7, b=5 and c=12
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-7 5 12
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = -1
x = 1.71428571
---
factorization:
-7xx + 5x + 12 = 0
xx - (5/7)x - 12/7 = 0
(x + 7/7)(x - 12/7)
(x + 1)(x - 1.71428571)
---
answer:
the longest dimension of the rectangle = 12 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php